8 1/2 multiple by 3 7/8 = estimated
Answer by Earlsdon(6294)

Estimate:
8 1/2 times 3 7/8 Notice that 3 7/8 is almost 4, so you can say:
8 1/2 times 4 = 4(8)+4(1/2) = 32 + 2 = 34 Estimated.